# SA-CAISR: Stage-Adaptive and Conflict-Aware Incremental Sequential Recommendation

### v1

@@ Abstract.
To overcome these challenges, we propose SA-CAISR, a Stage-Adaptive and Conflict-Aware Incremental Sequential Recommendation framework. As a buffer-free framework, SA-CAISR operates using only the old model and new data, directly addressing the high costs of replay-based techniques. SA-CAISR introduces a novel Fisher-weighted knowledge-screening mechanism that dynamically identifies outdated knowledge by estimating parameter-level conflicts between the old model and new data, allowing our approach to selectively remove obsolete knowledge while preserving compatible historical patterns. This dynamic balance between stability and adaptability allows our method to achieve a new state-of-the-art performance in incremental SR. Specifically, SA-CAISR improves Recall@20 by 2.0%, MRR@20 by 2.1%, and NDCG@20 by 2.1% on average across datasets, while reducing memory usage by 97.5% and training time by 46.9% compared to the best baselines. This efficiency allows real-world systems to rapidly update user profiles with minimal computational overhead, ensuring more timely and accurate recommendations. Our implementation can be found at: https://github.com/Togmm/SA-CAISR
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: TOISjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Information systems Recommender systemsccs: Computing methodologies Online learning settings
@@ 1. Introduction
In real-world applications, user behaviors are inherently non-stationary: new users and items continuously emerge, short-term trends coexist with long-term interests, and preferences drift over time (Zhang et al., 2019; Liu et al., 2024). This dynamic nature often leads to significant distribution shifts between historical and future interactions, challenging the model’s ability to generalize to out-of-distribution (OOD) samples (Zhang et al., 2024c). To address such shifts, SR models must be updated frequently. However, frequent retraining on the entire historical dataset is both time-consuming and computationally expensive, while naive incremental fine‑tuning on only the newly accumulated data typically leads to catastrophic forgetting of previously learned knowledge (Cai et al., 2022; Lopez-Paz and Ranzato, 2017; De Lange et al., 2022).
@@
To overcome these limitations, our objective is to design a buffer-free framework that eliminates the memory and privacy burdens of replay methods, while equipping the model with the ability to actively identify and filter conflicting knowledge—a capability missing in traditional regularization approaches. To this end, we propose the Stage-Adaptive and Conflict-Aware Incremental Sequential Recommendation (SA-CAISR) framework. As illustrated in Figure 1, unlike replay methods, SA-CAISR is buffer-free, thereby preserving privacy and operational efficiency. Distinct from traditional regularization, it introduces a Fisher-weighted knowledge-screening mechanism. This mechanism computes Fisher information on new data to quantify parameter-level conflicts, selectively masking parameters in the old model. This enables the model to dynamically discard outdated or conflicting knowledge while strictly preserving compatible historical patterns. To ensure the updated model retains access to valid historical signals, we further integrate an InfoNCE-based consistency loss. Experimental results validate the superior efficacy of SA-CAISR, demonstrating that it outperforms all baselines—including resource-intensive methods that utilize the complete historical data. Specifically, SA-CAISR improves Recall@20 by 2.0%, MRR@20 by 2.1%, and NDCG@20 by 2.1% on average across datasets. Beyond accuracy, the framework exhibits exceptional efficiency. While emerging lightweight architectures (Zhang et al., 2025) primarily focus on optimizing inference latency via structural designs, SA-CAISR tackles the computational bottleneck of continuous training. Consequently, it reduces memory usage by 97.5% and training latency by 46.9% compared to representative baselines, striking an optimal balance between high-performance recommendation and computational scalability.
@@
  • We propose a Fisher-weighted parameter screening mechanism that enables fine-grained filtering, allowing the model to effectively distinguish between valuable historical knowledge and conflicting or outdated information.
@@ 2. Related Work
Sequential recommendation (SR) aims to model users’ dynamic interests by leveraging their chronological interaction histories to predict the next item. Early SR research focused on traditional methods such as FPMC (Rendle et al., 2010). Later, neural models became dominant, including RNN-based approaches like GRU4Rec (Hidasi et al., 2016), Transformer-based methods such as SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019). To handle complex sequences and leverage external knowledge, efficient State Space Models (SSMs) (Yang et al., 2024) and LLM-based recommendation (Wang et al., 2025; Rajput et al., 2023) have emerged as rapidly growing directions. Notably, recent studies have extended these paradigms to generative retrieval and data-efficient tuning (Lin et al., 2024; Li et al., 2023), significantly enhancing representation capabilities but also increasing the computational cost of model updates. As user preferences and item pools evolve continuously in real-world applications, SR models must be updated frequently. This necessity gives rise to Incremental Sequential Recommendation (ISR), which focuses on enabling continual adaptation while mitigating catastrophic forgetting. Existing methods can be broadly categorized into two mainstream families: Replay-based approaches and Regularization-based approaches.
@@ 2.2. Regularization and Distillation Methods
Classic parameter-level regularization methods, such as Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017), aim to protect critical model parameters. EWC identifies these weights by evaluating the sensitivity of the model’s performance on data from previous stages to parameter updates, effectively penalizing changes to important weights. In the domain of recommendation, these ideas have been extended to preserve structural knowledge. GraphSAIL (Xu et al., 2020) enforces both local and global structural constraints on the embedding space to maintain historical user-item relationships. Similarly, SAIL-PIW (Wang et al., 2023) introduces personalized imitation weights, adaptively adjusting the regularization strength for each user based on their interaction frequency and preference volatility.
@@
While these methods avoid storing past data, a common limitation remains: they often treat the historical model (the “teacher”) as a fixed reference, constraining the new model to match it without explicitly modeling distribution shift or conflict between old and new knowledge (Farquhar and Gal, 2019). Our method, SA-CAISR, addresses this by integrating the conflict-awareness of parameter regularization with the semantic consistency of contrastive learning. Rather than distilling from a static teacher, we employ a Fisher-guided masking mechanism to dynamically filter out conflicting parameters from the reference model. This creates a ”purified” teacher for the InfoNCE loss, ensuring that the model preserves only compatible historical knowledge while freely adapting to new trends.
@@ 4.1.2. Adaptive Masking Mechanism
where $\gamma$ is a decay factor that gradually weakens the accumulated importance from previous stages (See Algorithm 1, step 20).
@@ 4.2. Consistency Preservation via Contrastive Alignment
Having screened out conflicting parameters, the next objective is to align the updated model with the valid historical knowledge preserved in the reference model. Instead of imposing rigid parameter-level constraints, we seek a flexible alignment in the representation space. Inspired by recent advances in soft contrastive learning that demonstrate superior robustness against sequence uncertainty (Zhang et al., 2024b), we employ the InfoNCE objective (Xie et al., 2022). The core logic is to maximize the similarity between the updated representation and the consistent historical representation, while differentiating them from other items.
@@
$\mathcal{L}_{\text{InfoNCE}}^{(k)}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\big(\mathrm{sim}(\mathbf{h}_{i},\mathbf{h}_{i}^{+})/\tau\big)}{\exp\big(\mathrm{sim}(\mathbf{h}_{i},\mathbf{h}_{i}^{+})/\tau\big)+\sum_{j\in\mathcal{N}^{(k)}_{i}}\exp\big(\mathrm{sim}(\mathbf{h}_{i},\mathbf{h}_{j})/\tau\big)}\,$,
@@ 4.3. Model Optimization
6:   (a) CE training:
7:   Optimize $W^{updated}$ by Eq. 18 to learn new knowledge
8:   Compute $\nabla_{W^{ref}}\mathcal{L}_{CE}$ on $B$ using Eq. 18
@@
16:   Compute the InfoNCE loss between $W^{updated}$ and $W^{ref}_{f}$ as defined in Eq. 17
@@
18:   (e) Final loss:
19:   Combine objectives and optimize $W^{updated}$ according to Eq. 19
20:  end for
21:  (f) Fisher decay:
22:  Apply decay to Fisher weights as specified in Eq. 16
23:end for
24:Result: $W_{t}=W^{updated}$
@@ 5.1.2. Baseline
To rigorously evaluate SA-CAISR, we compare it against several representative baselines widely used in incremental sequential recommendation, including replay-based methods (e.g., ADER) and regularization-based methods (e.g., EWC), as well as standard approaches such as fine-tuning, retraining, and linear models. These baselines cover a broad spectrum of strategies for retaining historical knowledge while learning from new data, providing a comprehensive context for assessing the effectiveness of our proposed framework. We deliberately restrict our study to attention-based sequential recommenders to control architectural confounders and isolate the effect of conflict-aware incremental learning: attention-only backbones are a dominant and standardized choice in modern SR, and focusing on them ensures that performance differences primarily reflect the incremental strategy rather than heterogeneous sequence encoders. Consequently, several recent methods—such as PISA (Yoo et al., 2025), SAIL-PIW (Wang et al., 2023), ReLoop2 (Zhu et al., 2023), and GraphSAIL (Xu et al., 2020)—are not included because they either depart from our attention-only setting or integrate additional architectural components beyond our scope. A further practical constraint is that these works do not provide publicly available implementations or sufficient experimental details, making faithful reproduction under consistent conditions challenging, especially for conflict-aware mechanisms. Instead, we focus on baselines for which reliable reproduction is feasible within the attention-based setting, enabling a fair and systematic assessment of SA-CAISR. This setup ensures that the reported improvements can be attributed to the Fisher-guided conflict detection and selective knowledge filtering framework, and provides a meaningful reference for future research in conflict-aware incremental sequential recommendation.
@@ 5.2. Overall Performance and Resource Efficiency Comparison (RQ1)
Table 2 and  3 summarize the overall recommendation performance, training time, and peak memory usage of all methods on four datasets. SA-CAISR consistently achieves the best performance across all metrics while maintaining efficiency comparable to lightweight regularization-based methods. On average, it improves Recall@20 by 2.0% over the strongest baseline while reducing memory usage by 97.5% and training time by 46.9% compared to the replay-based ADER. These results demonstrate that SA-CAISR effectively balances accuracy and efficiency, achieving robust long-term adaptation in dynamic recommendation scenarios.
@@ 5.3. Robustness under Varying Stage Intervals and Preference Shifts (RQ2)
Across all datasets, SA-CAISR demonstrates consistently strong robustness under different stage intervals and preference shift magnitudes. On average, it improves Recall@20 by 1.4–3.6%, MRR@20 by 0.8–3.2%, and NDCG@20 by 1.1–2.7% over the strongest baseline (ADER), highlighting its ability to effectively balance historical knowledge retention and adaptation to evolving user preferences.
@@ 7. Limitations and Future Work
While SA-CAISR achieves consistent gains with low overhead, our study has two categories of limitations. (i) Evaluation scope and comparability: we restrict backbones to attention-based sequential recommenders to isolate the effect of conflict-aware updating, which limits generality to graph- and state-space models; several recent ISR methods lack public code or sufficient detail, preventing strict comparisons under our setting. (ii) Methodological constraints: our conflict estimation operates at the parameter level and on a per‑stage basis, which does not explicitly capture user/item heterogeneity or disentangle short-term shocks from long-term drifts; stability–plasticity is governed by global hyperparameters (e.g., CE–InfoNCE weighting, fixed top‑K hard negatives), which may be suboptimal across users or stages.
In future work, we plan three focused extensions: (i) Personalize conflict modeling by augmenting Fisher-guided screening with user/item-level signals (e.g., uncertainty and gradient agreement) and separating short-term shocks from long-term drifts. (ii) Learn adaptive stability–plasticity control via per‑stage (or per‑user) weighting of CE and InfoNCE and dynamic hard‑negative selection, potentially with lightweight meta‑learning; (iii) Broaden architectural and system coverage by extending to graph-based and state-space backbones, exploring online/streaming settings, and benchmarking under public, reproducible pipelines to strengthen comparability.

### v2

@@ Abstract.
To overcome these challenges, we propose SA-CAISR, a Stage-Adaptive and Conflict-Aware Incremental Sequential Recommendation framework. As a buffer-free framework, SA-CAISR operates using only the old model and new data, directly addressing the high costs of replay-based techniques. SA-CAISR introduces a novel Fisher-weighted knowledge-screening mechanism that dynamically identifies outdated knowledge by estimating parameter-level conflicts between the old model and new data, selectively removing obsolete knowledge while preserving compatible historical patterns. This dynamic balance between stability and adaptability allows our method to achieve state-of-the-art performance in incremental SR. Specifically, SA-CAISR improves Recall@20 by 2.0% on average across datasets, while reducing memory usage by 97.5% and training time by 46.9% compared to the best baseline. This efficiency allows real-world systems to rapidly update user profiles with minimal computational overhead, ensuring more timely and accurate recommendations. Our implementation can be found at: https://github.com/Togmm/SA-CAISR
@@
$c$s: Information systems Recommender systemsccs: Computing methodologies Online learning settings
@@ 1. Introduction
In real-world applications, user behaviors are inherently non-stationary: new users and items continuously emerge, short-term trends coexist with long-term interests, and preferences drift over time (Zhang et al., 2019; Liu et al., 2024). Recent studies emphasize that efficiently capturing these shifts from continuous data streams is critical for maintaining recommendation accuracy (Lee et al., 2025b).This dynamic nature often leads to significant distribution shifts between historical and future interactions, challenging the model’s ability to generalize to out-of-distribution (OOD) samples (Zhang et al., 2024c). To address such shifts, SR models must be updated frequently. However, frequent retraining on the entire historical dataset is both time-consuming and computationally expensive, while naive incremental fine‑tuning on only the newly accumulated data typically leads to catastrophic forgetting of previously learned knowledge (Cai et al., 2022; Lopez-Paz and Ranzato, 2017; De Lange et al., 2022).
@@
To overcome these limitations, our objective is to design a buffer-free framework that eliminates the memory and privacy burdens of replay methods, while equipping the model with the ability to actively identify and filter conflicting knowledge—a capability missing in traditional regularization approaches. To this end, we propose the Stage-Adaptive and Conflict-Aware Incremental Sequential Recommendation (SA-CAISR) framework. As illustrated in Figure 1, unlike replay methods, SA-CAISR is buffer-free, thereby preserving privacy and operational efficiency. Distinct from traditional regularization, it introduces a Fisher-weighted knowledge-screening mechanism. This mechanism computes Fisher information on new data to quantify parameter-level conflicts, selectively masking parameters in the old model. This enables the model to dynamically discard outdated or conflicting knowledge while strictly preserving compatible historical patterns. To ensure the updated model retains access to valid historical signals, we further integrate an InfoNCE-based consistency loss. Experimental results validate the superior efficacy of SA-CAISR, demonstrating that it outperforms all baselines—including resource-intensive methods that utilize the complete historical data. Specifically, SA-CAISR improves Recall@20 by 2.0%, MRR@20 by 2.1%, and NDCG@20 by 2.1% on average across datasets. Beyond accuracy, the framework exhibits exceptional efficiency. While emerging lightweight architectures (Zhang et al., 2025b, a) primarily focus on optimizing inference latency via structural designs, SA-CAISR tackles the computational bottleneck of continuous training. Consequently, it reduces memory usage by 97.5% and training latency by 46.9% compared to representative baselines, striking an optimal balance between high-performance recommendation and computational scalability.
@@
  • We propose a framework that operates using only the old model and new data, eliminating the need for interaction replay buffers and significantly reducing resource overhead.
  • We introduce a novel mechanism that dynamically identifies and filters conflicting knowledge based on parameter-level importance, effectively balancing stability and plasticity.
@@ 2. Related Work
Sequential recommendation (SR) aims to model users’ dynamic interests by leveraging their chronological interaction histories to predict the next item. Early SR research focused on traditional methods such as FPMC (Rendle et al., 2010). Later, neural models became dominant, including RNN-based approaches like GRU4Rec (Hidasi et al., 2016), Transformer-based methods such as SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019). To handle complex sequences and leverage external knowledge, efficient State Space Models (SSMs) (Yang et al., 2024) and LLM-based recommendation (Wang et al., 2025a; Rajput et al., 2023) have emerged as rapidly growing directions. Notably, recent studies have extended these paradigms to generative retrieval and data-efficient tuning (Lin et al., 2024; Li et al., 2023), significantly enhancing representation capabilities but also increasing the computational cost of model updates. As user preferences and item pools evolve continuously in real-world applications, SR models must be updated frequently. This necessity gives rise to Incremental Sequential Recommendation (ISR), which focuses on enabling continual adaptation while mitigating catastrophic forgetting. Existing methods can be broadly categorized into two mainstream families: Replay-based approaches and Regularization-based approaches.
@@ 2.2. Regularization and Distillation Methods
Classic parameter-level regularization methods, such as Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017), aim to protect critical model parameters. EWC identifies these weights by evaluating the sensitivity of the model’s performance on data from previous stages to parameter updates, effectively penalizing changes to important weights. In the domain of recommendation, these ideas have been extended to preserve structural knowledge. GraphSAIL (Xu et al., 2020) enforces both local and global structural constraints on the embedding space to maintain historical user-item relationships. Similarly, SAIL-PIW (Wang et al., 2023b) introduces personalized imitation weights, adaptively adjusting the regularization strength for each user based on their interaction frequency and preference volatility.
@@
While these methods avoid storing past data, a common limitation remains: they often treat the historical model (the “teacher”) as a fixed reference, constraining the new model to match it without explicitly modeling distribution shift or conflict between old and new knowledge (Farquhar and Gal, 2019). Recent advances in sequential recommendation have shifted towards disentangling complex user intents to isolate valid patterns from noise or conflicting signals (Wang et al., 2025b). While such methods leverage Large Language Models (LLMs) to achieve high-quality disentanglement through semantic reasoning, they often incur substantial computational overhead unsuitable for real-time incremental updates. It is worth noting that a recent concurrent work, FGGM (Tan et al., 2026), also explores Fisher-guided gradient masking to address similar challenges in general continual learning. However, while sharing high-level similarities in leveraging Fisher Information, FGGM focuses on static classification tasks. In contrast, our SA-CAISR is specifically tailored for sequential recommendation, addressing the unique challenge of user preference drift through a stage-adaptive and conflict-aware mechanism. Our method, SA-CAISR, addresses this by integrating the conflict-awareness of parameter regularization with the semantic consistency of contrastive learning. Rather than distilling from a static teacher, we employ a Fisher-guided masking mechanism to dynamically filter out conflicting parameters from the reference model. This creates a ”purified” teacher for the InfoNCE loss, ensuring that the model preserves only compatible historical knowledge while freely adapting to new trends.
@@ 4.1.2. Adaptive Masking Mechanism
where $\gamma$ is a decay factor that gradually weakens the accumulated importance from previous stages (See Algorithm 1, step 21).
@@ 4.2. Consistency Preservation via Contrastive Alignment
Having screened out conflicting parameters, the next objective is to align the updated model with the valid historical knowledge preserved in the reference model. Instead of imposing rigid parameter-level constraints, we seek a flexible alignment in the representation space. Inspired by recent advances demonstrating that contrastive signals significantly enhance representation robustness against sequence uncertainty and noise (Wang et al., 2023a; Zhang et al., 2024b), we employ the InfoNCE objective (Xie et al., 2022). The core logic is to maximize the similarity between the updated representation and the consistent historical representation, while differentiating them from other items.
@@
(18) $\mathcal{L}_{\text{InfoNCE}}^{(k)}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\big(\mathrm{sim}(\mathbf{h}_{i},\mathbf{h}_{i}^{+})/\tau\big)}{\exp\big(\mathrm{sim}(\mathbf{h}_{i},\mathbf{h}_{i}^{+})/\tau\big)+\sum_{j\in\mathcal{N}^{(k)}_{i}}\exp\big(\mathrm{sim}(\mathbf{h}_{i},\mathbf{h}_{j})/\tau\big)}\,$,
@@ 4.3. Model Optimization
6:   (a) CE Loss Calculation:
7:   Compute $\mathcal{L}_{CE}$ using Eq. 19
8:   Compute gradients $\nabla_{W^{ref}}\mathcal{L}_{CE}$ on $B$ for Fisher calculation
@@
16:   Compute $\mathcal{L}_{InfoNCE}$ between $W^{updated}$ and $W^{ref}_{f}$ using Eq. 18
@@
18:   (e) Joint Optimization:
19:   $\mathcal{L}_{total}\leftarrow\mathcal{L}_{CE}+\alpha\cdot\mathcal{L}_{InfoNCE}$
20:   Update $W^{updated}$ by minimizing $\mathcal{L}_{total}$
21:  end for
22:  (f) Fisher decay:
23:  Apply decay to Fisher weights as specified in Eq. 16
24:end for
25:Result: $W_{t}=W^{updated}$
@@ 5.1.2. Baseline
To rigorously evaluate SA-CAISR, we compare it against several representative baselines widely used in incremental sequential recommendation, including replay-based methods (e.g., ADER) and regularization-based methods (e.g., EWC), as well as standard approaches such as fine-tuning, retraining, and linear models. These baselines cover a broad spectrum of strategies for retaining historical knowledge while learning from new data, providing a comprehensive context for assessing the effectiveness of our proposed framework. We deliberately restrict our study to attention-based sequential recommenders to control architectural confounders and isolate the effect of conflict-aware incremental learning: attention-only backbones are a dominant and standardized choice in modern SR, and focusing on them ensures that performance differences primarily reflect the incremental strategy rather than heterogeneous sequence encoders. Consequently, several recent methods—such as PISA (Yoo et al., 2025), SAIL-PIW (Wang et al., 2023b), ReLoop2 (Zhu et al., 2023), and GraphSAIL (Xu et al., 2020)—are not included because they either depart from our attention-only setting or integrate additional architectural components beyond our scope. A further practical constraint is that these works do not provide publicly available implementations or sufficient experimental details, making faithful reproduction under consistent conditions challenging, especially for conflict-aware mechanisms. Instead, we focus on baselines for which reliable reproduction is feasible within the attention-based setting, enabling a fair and systematic assessment of SA-CAISR. This setup ensures that the reported improvements can be attributed to the Fisher-guided conflict detection and selective knowledge filtering framework, and provides a meaningful reference for future research in conflict-aware incremental sequential recommendation.
@@ 5.2. Overall Performance and Resource Efficiency Comparison (RQ1)
Table 2 and  3 summarize the overall recommendation performance, training time, and peak memory usage of all methods on four datasets. SA-CAISR consistently achieves the best performance across all metrics while maintaining efficiency comparable to lightweight regularization-based methods. On average, it improves Recall@20 by 2.0% while reducing memory usage by 97.5% and training time by 46.9% compared to the strongest baseline(ADER). These results demonstrate that SA-CAISR effectively balances accuracy and efficiency, achieving robust long-term adaptation in dynamic recommendation scenarios.
@@ 5.3. Robustness under Varying Stage Intervals and Preference Shifts (RQ2)
Across all datasets, SA-CAISR demonstrates consistently strong robustness under different stage intervals and preference shift magnitudes. On average, it improves Recall@20 by 0.4–3.6%, MRR@20 by 0.8–1.8%, and NDCG@20 by 0.8–2.6% over the best-performing baselines across all datasets, highlighting its superior ability to effectively balance historical knowledge retention and adaptation to evolving user preferences.
@@ 7. Limitations and Future Work
While SA-CAISR establishes a robust buffer-free framework for incremental sequential recommendation, we identify specific limitations that contextualize our contributions and point toward future research directions.
Limitations
First, regarding architectural generalization, we restricted our evaluation to attention-based backbones (SASRec) to rigorously isolate the efficacy of our conflict-aware mechanism. We did not extend SA-CAISR to Graph Neural Networks (GNNs) or State Space Models (SSMs) in this study. This is due not only to the scarcity of reproducible implementations for recent baselines but also because adapting Fisher information estimation to graph topologies or recurrent states presents non-trivial theoretical challenges (e.g., structure-dependent conflicts) that warrant separate investigation. Second, our method relies on an independence assumption in conflict estimation to ensure linear computational complexity suitable for high-frequency updates. Our Fisher-weighted screening assumes a diagonal Fisher Information Matrix, ignoring covariance between weights. In deep networks where knowledge is entangled across neurons, this diagonal approximation may lead to suboptimal masking decisions compared to full-matrix methods. Third, concerning the granularity of conflict adaptation, while SA-CAISR is ”Stage-Adaptive” in capturing temporal distribution shifts, it remains ”User-Agnostic” within each stage. The conflict estimation operates at the global parameter level using stage-wide statistics, potentially struggling to disentangle short-term noise from long-term interest drifts for users with highly heterogeneous behavior patterns.
Future Work
To address these limitations, we propose three focused extensions. (1) Cross-Architecture Adaptation: We plan to generalize the framework to heterogeneous architectures by developing specialized Fisher estimation protocols for GNNs (handling structural forgetting) and SSMs (handling long-state forgetting). (2) High-Order Conflict Modeling: We aim to transcend the diagonal assumption by exploring efficient approximations of the full Fisher matrix, such as K-FAC. This would enable the detection of conflicts in the parameter correlation space, offering more precise protection of historical knowledge. (3) Instance-Aware Personalization: We intend to introduce fine-grained, instance-aware conflict detection. By incorporating user-specific uncertainty or gradient coherence signals, future versions could dynamically adjust the masking strength per user, effectively separating temporary behavioral shocks from genuine preference evolution.